# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ 3.1 Effective complexity of the model
We generalize previous results in a setting where the diversity condition may not hold. Starting from the conditional distribution captured by the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, we have that for every $y_{0}\in\mathcal{A}$,
@@
featuring that the conditional distribution ${p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})}$ is fully determined by the dot product ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$. Denote by ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ the orthogonal projectors onto $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, respectively. Since both ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ and ${\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$, the dot product between the two can be evaluated as:
@@
where we used the fact that ${\bm{\mathrm{P}}}_{\mathcal{F}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{F}}$ being an orthogonal projector. In general, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ may not commute [26]. We consider the subspaces:
@@
which will be central to our characterization. ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ underlie, respectively, the resulting projection of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ that contributes to the dot product. Only when ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, we have that $\mathcal{M}=\mathcal{N}=\mathrm{SIm}({\bm{\mathrm{f}}})\cap\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ [26]. In general, ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}\neq{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}\neq{\bm{\mathrm{g}}}_{0}$. To see this, consider the following example.
@@
Let $\{{\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2},{\bm{\mathrm{e}}}_{3}\}$ be an orthogonal basis of $\mathbb{R}^{3}$. Take $\mathcal{F}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2})$ and $\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{3})$. In this case, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, so it holds $\mathcal{M}=\mathcal{N}=\mathcal{F}\cap\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1})$.
@@
For two orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ and the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ defined on spaces of 7, it holds that (i) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}}))-\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}})-\mathrm{SIm}({\bm{\mathrm{g}}}_{0})^{\perp})$; (ii) $\mathcal{M}\subseteq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$; and (iii)
@@
All proofs can be found in Appendix B. As a consequence of this lemma, we can equivalently consider the projections ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ as the components of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ that are retained to evaluate the dot product of 8. This also means that whatever information is contained outside $\mathcal{M}$ and $\mathcal{N}$ is not relevant for the evaluation of the dot product. The dimensionality of $\mathcal{M}$ (and $\mathcal{N}$) can be viewed as a measure of model complexity since, intuitively, the larger $\dim(\mathcal{M})$ is, the more complex the relationship between $y$ and ${\bm{\mathrm{x}}}$ can be in $p(y\mid{\bm{\mathrm{x}}})$. For this reason, we call $\dim(\mathcal{M})$ the effective complexity of the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$.
Notice that, the effective complexity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is less than or equal to the representation dimensionality $d$.
@@ 3.2 Extended linear equivalence relation
Based on the effective complexity—which may be lower than the representation dimensionality—we introduce a weaker equivalence relation between models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$. Consider the spaces ~$\tilde{\mathcal{F}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{f}}})$ and $\tilde{\mathcal{G}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{g}}}_{0})$, and define ~$\tilde{\mathcal{M}}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})$ and $\tilde{\mathcal{N}}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})^{\perp}$.
@@
Two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ are extended-linearly equivalent, if both (1) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})$ and (2) there exist two full-rank matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{d\times\tilde{d}}$ defining, respectively, invertible transformations between $\mathcal{M}$ and $\tilde{}\mathcal{M}$, and between $\mathcal{N}$ and ~$\tilde{}\mathcal{N}$, such that ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$ and
@@
Figure 2: Illustration of the $\sim_{EL}$ equivalence relation. (Left) The model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ represents the embeddings over the manifold $\mathrm{Im}({\bm{\mathrm{f}}})\subsetneq\mathbb{R}^{3}$, yielding $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{3}$. To ease visualization, $\mathrm{Im}({\bm{\mathrm{f}}})$ is plotted as a continuous manifold in the figure, although this is not the case for actual textual inputs. The unembeddings lie on a two-dimensional space, $(\mathrm{SIm}({\bm{\mathrm{g}}}_{0}))\cong\mathbb{R}^{2}$, drawn in blue. Consequently the projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ map onto a two-dimensional subspace, i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{G}}$. (Right) The model $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ represents both the embeddings and the unembeddings on a two-dimensional space. We have that $\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}})=\mathrm{SIm}(\tilde{{\bm{\mathrm{g}}}_{0}})=\mathbb{R}^{2}$, which implies that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{I}}}$, thus applying these projection matrices does not modify the representations (top-right and bottom-right grids). (Center) The equivalence relation $\sim_{EL}$ specifies that both ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}$ as well as ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{{\bm{\mathrm{g}}}}$ are related by a linear invertible transformation given by the matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{3\times 2}$.
The above equivalence relation generalizes the linear equivalence already known in the literature [3, 20] to that of two models which can be related to each other on subspaces of dimension $\mathrm{dim}(\mathcal{M})\leq\min\{d,\tilde{d}\}$. It shows that, after projecting the representations to suitable equal-dimensional subspaces, namely $\mathcal{M},\mathcal{N},\tilde{}\mathcal{M}$, and ~$\tilde{}\mathcal{N}$, we can find an invertible linear transformation relating them. Here, the dimension of $\mathcal{M}$ and $\tilde{}\mathcal{M}$ are equal, requiring that two equivalent models share the same effective complexity. Furthermore, the condition between ${\bm{\mathrm{M}}}$ and ${\bm{\mathrm{N}}}$ preserves the dot-product:
@@
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)=\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$ (11)
As a consequence, models in the $\sim_{EL}$ equivalence class also satisfy $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. We show that it is an equivalence relation in Section B.3 and the explicit form of ${\bm{\mathrm{N}}}$ to make the definition hold. An example of this equivalence relation is illustrated in Figure 2.
@@ 3.3 Identifiability of next-token predictors
We can relate the representations of two models based on this equivalence relation when they express the same conditional distribution. The following theorem offers a characterization of models attaining the same conditional probability distribution:
@@
For all $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$, which may have different representation dimensions $d$ and $\tilde{d}$, we have
@@
With Theorem 5, we characterize the space of distributionally-equivalent next-token predictors. Models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ in the equivalence relation $\sim_{EL}$ are associated to a single conditional probability distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$. This is a one-to-one correspondence between the set of conditional probability distributions expressed by 1 and the set of equivalence classes of $\sim_{EL}$ (cf. Figure 1).
Interestingly, this characterization highlights the fact that our notion of effective complexity, which we defined for models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, is a well-defined complexity measure for distributions $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$ in the sense that it does not depend on the specific choice of parameterization. Indeed, the result implies that
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\implies\dim(\mathcal{M})=\dim(\tilde{}\mathcal{M})\,.$
Furthermore, We can show that the previous theorem reduces to linear identifiability when both the diversity condition and $d=\tilde{d}$ hold:
Corollary 6 (Adapted from Roeder et al. [3]).
@@
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})\ \text{and}\ {\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{M}}}^{-\top}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$

### v2

@@ 3.1 Effective complexity of the model
To generalize previous results to settings where the diversity condition may not hold, we introduce the notion of effective complexity of a model. Starting from the conditional distribution captured by the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, we have that for every $y_{0}\in\mathcal{A}$,
@@
indicating that the conditional distribution ${p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})}$ is fully determined by the dot product ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$. Denote by ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ the orthogonal projectors onto $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, respectively. Since both ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ and ${\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$, the dot product between the two can be evaluated as:
@@
where we used the fact that ${\bm{\mathrm{P}}}_{\mathcal{F}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{F}}$, as a property of orthogonal projectors. In general, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ may not commute [Rehder, 1980]. We consider the subspaces:
@@
which will be central to our characterization. Only when ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, we have that $\mathcal{M}=\mathcal{N}=\mathrm{SIm}({\bm{\mathrm{f}}})\cap\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ [Rehder, 1980]. In general, ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}\neq{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}\neq{\bm{\mathrm{g}}}_{0}$, as shown in the following example.
@@
Let $\{{\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2},{\bm{\mathrm{e}}}_{3}\}$ be an orthogonal basis of $\mathbb{R}^{3}$. Take $\mathcal{F}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2})$ and $\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{3})$. Then it follows that ${\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{e}}}_{1}{\bm{\mathrm{e}}}_{1}^{\top}+{\bm{\mathrm{e}}}_{2}{\bm{\mathrm{e}}}_{2}^{\top}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{e}}}_{2}{\bm{\mathrm{e}}}_{2}^{\top}+{\bm{\mathrm{e}}}_{3}{\bm{\mathrm{e}}}_{3}^{\top}$. In this case, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, so $\mathcal{M}=\mathcal{N}=\mathcal{F}\cap\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1})$.
@@
Given the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$, and the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ onto, respectively, the spaces $\mathcal{M}$ and $\mathcal{N}$, defined as in Equation (7), the following holds: (i) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}}))-\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}})-\mathrm{SIm}({\bm{\mathrm{g}}}_{0})^{\perp})$; (ii) $\mathcal{M}\subseteq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$; and (iii)
@@
All proofs can be found in Appendix B. As a consequence of this lemma, we can view the projections ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ as the parts of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ that are effectively retained when evaluating the dot product on the left-hand side of Equation (8). This also means the dot product depends solely on the projection of the embeddings onto $\mathcal{M}$ and the unembeddings onto $\mathcal{N}$; components of the embeddings and unembeddings which are orthogonal to these subspaces do not contribute. The dimensionality of $\mathcal{M}$ (and $\mathcal{N}$) can be viewed as a measure of model complexity since, intuitively, the larger $\dim(\mathcal{M})$ is, the more expressive the resulting model, and thus the more complex the relationship between $y$ and ${\bm{\mathrm{x}}}$, captured by $p(y\mid{\bm{\mathrm{x}}})$, can be. For this reason, we call $\dim(\mathcal{M})$ the effective complexity of the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$. Note that the effective complexity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is less than or equal to the representation dimensionality $d$.
@@ 3.2 Extended linear equivalence relation
We now introduce an equivalence relation among models with potentially different representation dimensionality, building on our notion of effective complexity. We consider next-token prediction models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, and spaces ~$\tilde{\mathcal{F}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{f}}})$, $\tilde{\mathcal{G}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{g}}}_{0})$, ~$\tilde{\mathcal{M}}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})$ and $\tilde{\mathcal{N}}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})^{\perp}$, as introduced in Section 3.1.
@@
Two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ are extended-linearly equivalent, if both (i) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})$ and (ii) there exist two full-rank matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{d\times\tilde{d}}$ defining, respectively, invertible transformations from $\mathcal{M}$ to $\tilde{}\mathcal{M}$, and from $\mathcal{N}$ to ~$\tilde{}\mathcal{N}$, such that ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$ and
@@
The above equivalence relation generalizes the linear equivalence already known in the literature [Roeder et al., 2021, Khemakhem et al., 2020a] to that of two models which can be related to each other on subspaces of dimension ${\mathrm{dim}(\mathcal{M})\leq\min\{d,\tilde{d}\}}$. It shows that, after projecting the representations to suitable equal-dimensional subspaces, namely $\mathcal{M},\mathcal{N},\tilde{}\mathcal{M}$, and ~$\tilde{}\mathcal{N}$, we can find an invertible linear transformation relating them. Here, the dimensions of $\mathcal{M}$ and $\tilde{}\mathcal{M}$ are equal, requiring that two equivalent models share the same effective complexity. Furthermore, models that are $\sim_{EL}$-equivalent encode the same dot-product:
@@
As a consequence, models in the $\sim_{EL}$ equivalence class also satisfy $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. In Section B.3, we prove that Definition 3 is an equivalence relation and we provide the explicit form of the matrix ${\bm{\mathrm{N}}}$. The extended linear equivalence relation is illustrated in Figure 2.
@@ 3.3 Identifiability of next-token predictors
The following theorem provides a characterization of models generating the same conditional probability distribution (i.e., distributionally-equivalent next-token predictors):
@@
For all $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$, with representation dimensions $d$ and $\tilde{d}$ (not necessarily equal),
@@
In words, there is a one-to-one correspondence between the set of conditional probability distributions expressed in Equation (1) and the set of equivalence classes entailed by $\sim_{EL}$ (cf. Figure 1): models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ which are $\sim_{EL}$-equivalent can be mapped to a single conditional probability distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$.
Interestingly, Theorem 5 highlights the fact that our notion of effective complexity, defined in Section 3.1 for next-token prediction models, is a well-defined complexity measure for a distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$, in the sense that it does not depend on the specific choice of embedding and unembedding functions $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$. Indeed, the result implies that
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\implies\dim(\mathcal{M})=\dim(\tilde{}\mathcal{M})\,.$ (13)
Furthermore, as a special case of Theorem 5, when the diversity condition holds and $d=\tilde{d}$ we recover known results on linear identifiability:
Corollary 6 (Adapted from [Roeder et al., 2021]).
@@
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})\ \text{and}\ {\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{M}}}^{-\top}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$ (15)
In Section B.5, we provide an example about non-linear distortions that can arise in models that are $\sim_{EL}$-equivalent but not diverse (Definition 1).
3.4 Implications for empirical practice